# Optimal Computation in Leaderless and Multi-Leader Disconnected Anonymous Dynamic Networks

### 20 Feb 2024 v5

@@ 32   Return $(n^{\prime},t)$
It remains to prove that $n^{\prime}\neq-1$. Since $\tau$ exists by assumption, the error code $-1$ cannot be returned at Line 6, and can only be returned at Line 30. In turn, this can only occur if a counting cut has been found and $\ell^{\prime}<\ell$. However, we have already proved that no level up to $L_{t}$ is missing any nodes, which implies that the counting cut contains nodes representing all processes, and in particular $\ell^{\prime}=\delta\ell$. Since $\delta\geq 1$, we have $\ell^{\prime}\geq\ell$, and the condition at Line 30 is not satisfied. ∎
Appendix B Survey of Related Work
We examine related work on Counting and Average Consensus, distinguishing between the case of dynamic networks with unique IDs, the case of static anonymous networks, and the case of dynamic anonymous networks.
B.1 Dynamic Networks with IDs
The problem of counting the size of a dynamic network has been first studied by the peer-to-peer systems community [39]. In this case having an exact count of the network at a given time is impossible, as processes may join or leave in an unrestricted way. Therefore, their algorithms mainly focus on providing estimates on the network size with some guarantees. The most related is the work that introduced $1$-union-connected networks [36]. They show a counting algorithm that terminates in at most $n+1$ rounds when messages are unrestricted and in $O(n^{2})$ rounds when the message size is $O(\log n)$ bits. The techniques used heavily rely on the presence of unique IDs and cannot be extended to our settings.
B.2 Anonymous Static Networks
The study of computability on anonymous networks has been pioneered by Angluin in [1] and it has been a fruitful research topics for the last 30 years [1, 7, 12, 13, 14, 27, 49, 52]. A key concept in anonymous networks is the symmetry of the system; informally, it is the indistinguishability of processes that have the same view of the network. As an example, in an anonymous static ring topology, all processes will have the exact same view of the system, and such a view does not change between rings of different size. Therefore, non-trivial computations including counting are impossible on rings, and some symmetry-breaking assumption is needed (such as a leader [27]). The situation changes if we consider topologies that are asymmetric. As an example, on a wheel graph the central process has a view that is unique, and this allows for the election of a leader and the possibility, among other tasks, of counting the size of the network.
Several tools have been developed to characterize what can be computed on a given network topology (examples are views [52] or fibrations [8]). Unfortunately, these techniques are usable only in the static case and are not defined for highly dynamic systems like the ones studied in our work. Regarding the counting problem in anonymous static networks with a leader, [40] gives a counting algorithm that terminates in at most $2n$ rounds.
B.3 Counting in Anonymous Interval-Connected Networks
The papers that studied counting in anonymous dynamic networks can be divided into two periods. A first series of works [11, 20, 21, 40] gave solutions for the counting problem assuming some initial knowledge on the possible degree of a process. As a matter of fact [40] conjectured that some kind of knowledge was necessary to have a terminating counting algorithm. A second series of works [18, 24, 29, 31, 32, 34, 33] has first shown that counting was possible without such knowledge, and then has proposed increasingly faster solutions, culminating with the linear time asymptotically optimal solution of [24]. We remark that all these papers assume that a leader (or multiple leaders in [32]) is present. This assumption is needed to deterministically break the system’s symmetry.
Counting with knowledge on the degrees.
Counting in interval-connected anonymous networks was first studied in [40], where it is observed that a leader is necessary to solve counting in static (and therefore also dynamic) anonymous networks (this result can be derived from previous works on static networks such as [8, 52]). The paper does not give a counting algorithm but it gives an algorithm that is able to compute an upper bound on the network size. Specifically, [40] proposes an algorithm that, using an upper bound $d$ on the maximum degree that each process will ever have in the network, calculates an upper bound $U$ on the size of the network; this upper bound may be exponential in the actual network size ($U\leq d^{n}$).
Assuming the knowledge of an upper bound on the degree, [20] given a counting algorithm that computes $n$. Such an algorithm is really costly in terms of rounds; it has been shown in [11] to be doubly exponential in the network size. The algorithm proposes a mass distribution approach akin to local averaging [51].
An experimental evaluation of the algorithm in [20] can be found in [22]. The result of [20] has been improved in [11], where, again assuming knowledge of an upper bound $d$ on the maximum degree of a process, an algorithm is given that terminates in $O\left(n(2d)^{n+1}\frac{\log n}{\log{d}}\right)$ rounds. A later paper [21] has shown that counting is possible when each process knows its degree before starting the round (for example, by means of an oracle). In this case, no prior global upper bound on the degree of processes is needed. [21] only show that the algorithm eventually terminates but does not bound the termination time.
We remark that all the above works assume some knowledge on the dynamic network, as an upper bound on the possible degrees, or as a local oracle. Moreover, all of these works give exponential-time algorithms.
Counting without knowledge on the degrees.
The first work proposing an algorithm that does not require any knowledge of the network was [19]. The paper proposed an exponential-time algorithm that terminates in $O\left(n^{n+4}\right)$ rounds. Moreover, it also gives an asymptotically optimal algorithm for a particular category of networks (called persistent-distance). In this type of network, a process never changes its distance from the leader.
This result was improved in [29, 31], which presented a polynomial-time counting algorithm. The paper proposes Methodical Counting, an algorithm that counts in $O(n^{5}\log^{2}(n))$ rounds. Similarly to [20, 21], the paper uses a mass-distribution process that is coupled with a refined analysis of convergence time and clever techniques to detect termination. The paper also notes that, using the same algorithm, all algebraic and boolean functions that depend on the initial number of processes in a given state can be computed. In [30, 34], the same authors extended their result to networks where $\ell\geq 1$ leaders are present (with $\ell$ known in advance), and gave an algorithm that terminates in $O\left(\frac{n^{4+\epsilon}}{\ell}\log^{3}(n)\right)$ rounds for any $\epsilon>0$. In particular, when $\ell=1$, this result improves on the running time of [29, 31]. Improve the running time and deriving tight bounds for counting in the multi-leader case was left as an open problem in [30, 34]. We give a definitive answer to these questions by providing a better running time that is optimal for any constant number of leaders.
Finally, in [32], they show a counting algorithm parameterized by the isoperimetric number of the dynamic network. The technique used is similar to [29, 31], and it uses the knowledge of the isoperimetric number to shorten the termination time. Specifically, for adversarial graphs (i.e., with non-random topology) with $\ell$ leaders ($\ell$ is assumed to be known in advance), they give an algorithm terminating in $O\left(\frac{n^{3+\epsilon}}{\ell{i_{min}}^{2}}\log^{3}(n)\right)$ rounds, where $i_{min}$ is a known lower bound on the isoperimetric number of the network. This improves the work in [34], but only in graphs where $i_{min}$ is $\omega({1}/{\sqrt{n}})$. The authors also study various types of graphs with stochastic dynamism; we remark that in this case they always obtain superlinear results, as well. The best case is that of Erdős–Rényi, graphs where their algorithm terminates in $O\left(\frac{n^{1+\epsilon}}{\ell{p_{min}}^{2}}\log^{5}(n)\right)$ rounds; here $p_{min}$ is the smallest among the probabilities of creating an arc on all rounds. Specifically, if $p_{min}=O({1}/{n})$, their algorithm is at least cubic.
A recent breakthrough has been shown in [24], which proposed the novel technique of history trees. A history tree is a combinatorial structure that models the entire evolution of an anonymous dynamic graph. By developing a theory of history trees for dynamic networks with a unique leader, the authors have shown a terminating solution for the Generalized Counting problem in $33n-$ 3 rounds. [^20] In the Generalized Counting problem, each process starts with a certain input, and the problem is solved when each process has computed the multiset of these inputs. The authors have shown that the Generalized Counting problem is complete for the class of problems solvable in general dynamic networks, proving that computing in anonymous dynamic networks with a unique leader is linear. The authors have also given a stabilizing, non-terminating, algorithm for Generalized Counting that stabilizes in roughly $2n$ rounds, providing an almost matching lower bound (we will discuss lower bounds below). We remark that the results of [24] do not apply to the leaderless or the multi-leader cases; the latter is left as an open problem in [24], which we settle in the present paper.
All the above works assume the dynamic network to be connected at each round. The only work that studied counting in disconnected networks is the recent pre-print [33]. The paper proposes an algorithm that solves the Generalized Counting in $\tilde{O}\left((R+\frac{n}{\ell})\frac{n^{2T(1+\epsilon)}}{i^{2}_{min}}\right)$ rounds, [^21] A function is $\tilde{O}(f(n))$ if it is $O(f(n)g(n))$ for some polylogarithmic function $g(n)$. where $T$ is the dynamic disconnectivity, $\ell$ is the number of leaders, $i_{min}$ is the isoperimetric number of the network, and $R$ is the cumulative bit length of all inputs. In unknown networks, assuming inputs of small size, the number of rounds is roughly $\tilde{O}\left(\frac{n^{2T(1+\epsilon)+3}}{\ell}\right)$. Interestingly, the algorithm requires messages of size $O(\log n)$ bits. We highlight that the algorithm has a running time of roughly $\tilde{O}\left(\frac{n^{5}}{\ell}\right)$ in dynamic networks that are connected at each round, and the complexity grows exponentially with the dynamic disconnectivity $T$.
Summarizing, to the best of our knowledge, there exists no worst-case cubic-time algorithm for the Generalized Counting with multiple leaders and no algorithm that scales linearly with the dynamic disconnectivity and the size of the network.
Lower bounds on counting.
From [36], a trivial lower bound of $n-1$ rounds can be derived, as counting obviously requires information from each process to be spread in the network. The first non-trivial lower bound for general dynamic networks has been given in the recent work [24]: any algorithm that stabilizes on the correct count requires $62n-$ 6 rounds. We remark that such a lower bound also holds for terminating algorithms. Another interesting lower bound is in [18], which shows a specific category of anonymous dynamic networks with constant temporal diameter (the time needed to spread information from a process to all others is at most 3 rounds), but where counting requires $\Omega(\log n)$ rounds.
B.4 Average Consensus
In the Average Consensus problem, each process $v_{i}$ starts with an input value $x_{i}(0)$, and the goal is to compute the average of these initial values. This problem has been studied for decades in the communities of distributed control and distributed computing [6, 15, 16, 17, 32, 42, 45, 47, 51, 54]. In the following, we give an overview that places our result in the current body of knowledge. A more detailed picture can be found in the surveys [26, 43, 46]. We can divide current papers between the ones that give convergent solutions to Average Consensus and the ones that give finite-time solutions.
B.5 Convergent Average Consensus
In convergent Average Consensus algorithms, the consensus is not reached in finite time, but each process has a local value that asymptotically converges to the average. A prototypical family of solutions [6, 51] is based on the so-called convex combination algorithms, where each process updates its local value $x_{i}(r)$ at every round $r$ as follows:
$x_{i}(r)=\sum_{\forall{v_{j}\in N(r,v_{i})\cup\{v_{i}\}}}a_{ij}(r)\cdot x_{j}(r-1).$
The value $a_{ij}(r)$ is taken from a weight matrix that models a dynamic graph. We remark that convex combination algorithms do not need unique IDs, and thus work in anonymous networks.
The $\epsilon$-convergence of an algorithm is defined as the time it takes to be sure that the maximum discrepancy between the local value of a process and the mean is at most $\epsilon$ times the initial discrepancy. That is, if the mean is $m=\sum_{i\in V}x_{i}(0)/{|V|}$, the following should hold:
$\frac{\text{max}_{i}\left\{|x_{i}(r)-m|\right\}}{\text{max}_{i}\left\{|x_{i}(0)-m|\right\}}\leq\epsilon.$
The local averaging approach has been studied in depth, and several upper and lower bounds for $\epsilon$-convergence are known for both static and dynamic networks [46, 47]. The procedure $\epsilon$-converges in $O\left(Tn^{3}\log(\frac{1}{\epsilon})\right)$ rounds if, at every round, the weight matrix $A(r)$ such that $(A(r))_{ij}=a_{ij}(r)$ is doubly stochastic (i.e., the sum of the values on rows and columns is $1$) and the dynamic graph is $T$-interval-connected [42]. In a dynamic network, it is possible to have doubly stochastic weight matrices when an upper bound on the processes’ degrees is known [17]. Without such knowledge, if the dynamic graph is always connected and stable (i.e., it changes every two rounds), then it is possible to implement a Metropolis weights strategy that converges in $O\left(n^{2}\log(\frac{n}{\epsilon})\right)$ rounds [42].
Numerous other studies have investigated averaging algorithms based on Metropolis rules. However, Metropolis rules require processes to know their out-degree prior to the broadcast phase of each round, making them unsuitable for our model. Charron-Bost et al. observed this in [16]: “Unfortunately, local algorithms cannot implement the Metropolis rule over dynamic networks. The rule is only “local” in the weak sense that an agent’s next estimate $x_{i}(t)$ depends on information present within distance $2$ of agent $i$ in the communication graph $\mathbb{G}(t)$, which is not local enough when the network is subject to change”. The only paper that assumes a similar setting to ours is [16], but it restricts the dynamic graph to be $1$-interval-connected. The paper shows an algorithm that uses MaxMetropolis weights and converges in $O\left(n^{4}\log(\frac{n}{\epsilon})\right)$ rounds.
We remark that all of the above works only achieve convergence to the average of the inputs, and do not stabilize on the average in finite time.
B.6 Finite-Time Average Consensus
The algorithms in the second class solve the finite-time Average Consensus; in this case, the value stabilizes to the actual average in a finite number of rounds. The majority of the literature on finite-time Average Consensus has considered static networks [26, 45, 54]. In such a setting, algorithms that stabilize in a linear number of rounds are known [45, 54]. Few works considered anonymous dynamic networks; [42] describes an algorithm that stabilizes in $O(n^{2})$ rounds and requires the network to change every three rounds, while in [15] a randomized Monte Carlo linear algorithm is given. An interesting take is given in [32], which investigates terminating Average Consensus algorithms for adversarial dynamic graphs and random dynamic graphs (i.e., Watts–Strogatz, Barabàsi–Albert, RGG, and Erdős–Rényi–Gilbert graphs). The algorithm for the adversarial case has a time complexity of $O\left(\frac{n^{5}}{\ell}\log^{3}({n})\right)$ rounds, where $\ell$ is the (known) number of leaders in the system. Our multi-leader algorithm is optimal for any constant number of leaders, and settles an open problem of [32] on finding tight results. For random dynamic graphs, the complexity of the algorithm changes according to the model, but all the algorithms presented in [32] are super-linear and require the knowledge of the number of leaders to terminate.
To the best of our knowledge, there is no deterministic solution to Average Consensus that stabilizes, or even converges, in a linear number of rounds in unknown dynamic networks.

### 05 Apr 2024 v6

@@ 32   Else return $(n^{\prime},t)$
It remains to prove that $n^{\prime}\neq-1$. Since $\vartheta$ exists by assumption, the error code $-1$ cannot be returned at Line 6 and can only be returned at Line 30. In turn, this can only occur if a counting cut has been found and $\ell^{\prime}<\ell$. However, we have already proved that no level up to $L_{t}$ is missing any nodes, which implies that the counting cut contains nodes representing all processes, and in particular $\ell^{\prime}=\delta\ell$. Since $\delta\geq 1$, we have $\ell^{\prime}\geq\ell$, and the condition at Line 30 is not satisfied. ∎
Worst-case example. For $\ell=\tau=1$, our Counting algorithm in Theorem 5.2 yields a running time of $23n-$ 2 rounds. The example in Figure 8, which can easily be generalized to networks of any size $n$, shows that our Counting algorithm can in fact terminate in $33n-$ 3 rounds. Indeed, the last node in a counting cut is at level $L_{2n-3}$, and then it takes an extra $n$ rounds for the terminating condition $t\geq t^{\ast}+n^{\ast}$ of Listing 2 to be satisfied.
Figure 8: An example of a dynamic network with $\ell=\tau=1$ and $n=6$ where the algorithm of Theorem 5.2 terminates in $33n-$ 3 rounds, almost matching the upper bound of $23n-$ 2 rounds. The white nodes in the history tree are not in the view of the leader at the last round. The numbers inside nodes represent guesses; the blue ones correspond to nodes that are marked as counted. The six highlighted nodes constitute a counting cut.
6 Negative Results
In this section we collect several negative results and counterexamples, which not only provide lower bounds asymptotically matching our algorithms’ running times, but also justify all of the assumptions made in Sections 4 and 5 about the a-priori knowledge of processes. Although some of these facts were previously known, here we offer simple and self-contained proofs based on history trees.
6.1 Unsolvable Problems
Showing that certain problems are not solvable in certain network models allows us to argue that our algorithms are universal, i.e., they can be applied to all solvable problems. Remarkably, all of our algorithms exhibit the highest level of generality possible.
Networks with leaders. We will prove that the multiset-based functions introduced in Section 2 are the only functions that can be computed deterministically in anonymous networks with leaders. This result, together with Proposition 2.1, shows that the algorithms in Sections 4.4 and 5 can be generalized to all functions that can be computed in networks with leaders, i.e., the multiset-based functions.
Proposition 6.1.
No function other than the multiset-based functions can be computed (with or without termination), even when restricted to connected static simple networks with a known and arbitrary number of leaders.
Proof.
Let us consider the (static) network whose topology at round $t$ is the complete graph $G_{t}=K_{n}$, i.e., each process receives messages from all other processes at every round. We can prove by induction that all nodes of the history tree other than the root have exactly one child. This is because any two processes with the same input always receive equal multisets of messages, and are therefore always indistinguishable. Thus, the history tree is completely determined by the multiset $\mu_{\lambda}$ of all processes’ inputs; moreover, a process’ view at any given round only depends on the process’ own input and on $\mu_{\lambda}$. By the fundamental theorem of history trees Theorem 3.1, this is enough to conclude that if a process’ output stabilizes, that output must be a function of the process’ own input and of $\mu_{\lambda}$, which is the defining condition of a multiset-based function. ∎
The following result justifies the assumption made in Sections 4.4 and 5 that processes have a-priori knowledge of the number of leaders $\ell$ in the system.
Proposition 6.2.
No algorithm can compute the Counting function $F_{C}$ (with or without termination) with no knowledge about $\ell$, even when restricted to connected static simple networks with a known and arbitrarily small ratio $\ell/n$.
Proof.
Leaderless networks. We will now prove that the frequency-based functions introduced in Section 2 are the only functions that can be computed deterministically in anonymous networks without leaders. This result, together with Proposition 2.2, shows that the algorithms in Sections 4.2 and 4.3 can be generalized to all functions that can be computed in leaderless networks, i.e., the frequency-based functions.
Although the next theorem is implied by [30, Theorem III.1], here we provide an alternative and simpler proof.
Proposition 6.3.
No function other than the frequency-based functions can be computed with no leader, even when restricted to connected static simple networks.
Proof.
Let $m_{1}$, $m_{2}$, …, $m_{k}$ be integers greater than $2$ with $\gcd(m_{1},m_{2},\dots,m_{k})=1$, and let $B$ be the complete $k$-partite graph with partite sets $V_{1}$, $V_{2}$, …, $V_{k}$ of sizes $m_{1}$, $m_{2}$, …, $m_{k}$, respectively. For any positive integer $\alpha$, construct the static network $G_{\alpha}$ consisting of $\alpha$ disjoint copies of $B$, augmented with $k$ cycles $C_{1}$, $C_{2}$, …, $C_{k}$ such that, for each $1\leq i\leq k$, the cycle $C_{i}$ spans all the $\alpha m_{i}$ processes in the $\alpha$ copies of $V_{i}$. Clearly, $G_{\alpha}$ is a connected static simple network.
Due to the fundamental theorem of history trees Theorem 3.1, all the processes with input $z_{i}$ must give the same output $\psi(z_{i},\mu_{\lambda})=\psi(z_{i},\mu_{\lambda_{\alpha}})=\psi(z_{i},\alpha\cdot\mu_{\lambda})$, regardless of $\alpha$. Hence, by definition, only frequency-based functions can be computed in these networks. ∎
The following result justifies the assumptions made in Section 4.3 that processes have knowledge of an upper bound on $n$ or on the dynamic diameter $d$ of the network.
Proposition 6.4.
No algorithm can solve the leaderless Average Consensus problem with explicit termination if nothing is known about the size of the network or its diameter, even when restricted to connected static simple networks.
Proof.
6.2 Lower Bounds
We will now give some lower bounds on the complexity of problems for anonymous dynamic networks. Since our algorithms have linear running times, our focus is on optimizing the multiplicative constants of the leading terms.
Preliminary results. We first prove some simple statements that will be used to derive lower bounds for stabilizing and terminating algorithms.
Lemma 6.5.
Let $\mathcal{G}$ and $\mathcal{G}^{\prime}$ be two networks on $n$ and $n^{\prime}$ processes respectively, where $n\neq n^{\prime}$. Assume that there is a process $p$ in $\mathcal{G}$ and a process $p^{\prime}$ in $\mathcal{G}^{\prime}$ such that $p$ and $p^{\prime}$ have isomorphic views at round $t$. Then,
  • No algorithm computes the Counting function $F_{C}$ and stabilizes within $t$ rounds in both $\mathcal{G}$ and $\mathcal{G}^{\prime}$.
  • No algorithm computes the Counting function $F_{C}$ in both $\mathcal{G}$ and $\mathcal{G}^{\prime}$ and terminates within $t$ rounds in $\mathcal{G}$ (or $\mathcal{G}^{\prime}$).
Proof.
Since $p$ and $p^{\prime}$ have isomorphic views at round $t$, they have isomorphic views at all rounds up to $t$. Thus, by Theorem 3.1, if $p$ and $p^{\prime}$ execute the same algorithm, they give equal outputs up to round $t$. Since the Counting function $F_{C}$ prescribes that $p$ must output $n$ and $p^{\prime}$ must output $n^{\prime}\neq n$, it is impossible for both processes to simultaneously give the correct output within $t$ rounds. In particular, no Counting algorithm can stabilize within $t$ rounds.
Moreover, if the execution of $p$ terminated within $t$ rounds, then the execution of $p^{\prime}$ would terminate at the same round, as well (again, due to Theorem 3.1). In that case, both processes would return the same output, which would be incorrect for at least one of them. In particular, no Counting algorithm can terminate within $t$ rounds in $\mathcal{G}$. ∎
Corollary 6.6.
Let $(\mathcal{G}_{n})_{n\geq 1}$ be a sequence where $\mathcal{G}_{n}$ is a network on $n$ processes. Assume that there exists a constant $k>0$ such that, for infinitely many values of $n\in\mathbb{N}$, there is a process in $\mathcal{G}_{n}$ and a process in $\mathcal{G}_{n+k}$ that have isomorphic views at round $an+b$, for integer constants $a>0$ and $b$. Then, there is no algorithm that computes the Counting function $F_{C}$ in every $\mathcal{G}_{n}$ in less than $an-ak+b+1$ rounds. If termination is required, the bound improves to $an+b+1$ rounds.
Proof.
Let $m\in\mathbb{N}$ be such that $\mathcal{G}_{m}$ and $\mathcal{G}_{m+k}$ contain processes with isomorphic views at round $t=am+b$. Assume for a contradiction that a Counting algorithm stabilized in $f(n)\leq an-ak+b$ rounds in every $\mathcal{G}_{n}$. Then, it would stabilize in $f(m)\leq am-ak+b\leq t$ rounds in $\mathcal{G}_{m}$ and in $f(m+k)\leq a(m+k)-ak+b=t$ rounds in $\mathcal{G}_{m+k}$. This contradicts Lemma 6.5, which states that no Counting algorithm stabilizes within $t$ rounds in both networks.
Moreover, if a Counting algorithm terminated in $f(n)\leq an+b$ rounds in every $\mathcal{G}_{n}$, it would terminate in $f(m)\leq am+b=t$ rounds in $\mathcal{G}_{n}$, which again contradicts Lemma 6.5. ∎
Unique-leader networks. We will now prove a lower bound of roughly $2n$ rounds on the Counting problem for always connected networks with a unique leader.
We first introduce a family of 1-union-connected dynamic networks. For any $n\geq 1$, we consider the dynamic network $\mathcal{G}_{n}$ whose topology at round $t$ is the graph $G^{(n)}_{t}$ defined on the system $\{p_{1},p_{2},\dots,p_{n}\}$ as follows. If $t\geq n-2$, then $G^{(n)}_{t}$ is the path graph $P_{n}$ spanning all processes $p_{1}$, $p_{2}$, …, $p_{n}$ in order. If $31\leq t\leq n-$ 3, then $G^{(n)}_{t}$ is $P_{n}$ with the addition of the single edge $\{p_{t+1},p_{n}\}$. We assume $p_{1}$ to be the leader and all other processes to have the same input.
Proposition 6.7.
In 1-union-connected dynamic simple networks with a unique leader, no algorithm can compute the Counting function $F_{C}$ in less than $62n-$ 6 rounds. If termination is required, the bound improves to $42n-$ 4 rounds.
Proof.
Let us consider the network $\mathcal{G}_{n}$ as defined above. It is straightforward to prove by induction that, at every round $t\leq n-3$, the process $p_{t+1}$ gets disambiguated, while all processes $p_{t+2}$, $p_{t+3}$, …, $p_{n}$ are still indistinguishable. So, the history tree of $\mathcal{G}_{n}$ has a very regular structure, which is illustrated in Figures 9 and 10. By comparing the history trees of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$, we see that the leaders of the two systems have identical views up to round $52n-$ 5. Our claim now follows immediately from Corollary 6.6, with $k=1$, $a=2$, and $b=-5$
A slightly better lower bound can be obtained if we allow self-loops in the network. Consider the static network $\mathcal{G}_{n}$ consisting of a path graph spanning all $n$ processes, where one endpoint of the path is the unique leader and the other endpoint has a self-loop. It is easy to see that the leaders in $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ have identical views up to round $22n-$ 2. This implies a lower bound of $32n-$ 3 rounds for stabilization and $12n-$ 1 rounds for termination in static networks with self-loops and a unique leader.
Figure 9: The first rounds of the dynamic network $\mathcal{G}_{n}$ used in Proposition 6.7 (left) and the corresponding levels of its history tree (right), where $n=6$; the process in blue is the leader. The white nodes and the dashed edges in the history tree are not in the view of the leader at round $7$. The labels $p_{1}$, …, $p_{6}$ have been added for the reader’s convenience, and mark the processes that get disambiguated, as well as their corresponding nodes of the history tree, which have anonymity $1$.
Figure 10: The first rounds of the dynamic network $\mathcal{G}_{n+1}$ with $n=6$. Observe that the view of the leader at round $7$ is identical to the view highlighted in Figure 9. The intuitive reason is that, from round $1$ to round $n-3$, both networks have a cycle whose processes are all indistinguishable (and are therefore represented by a single node in the history tree), except for the one process with degree $3$. Thus, the history trees of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ are identical up to level $n-3$. After that, the two networks get disambiguated, but this information takes another $n-3$ rounds to reach the leader. Therefore, if the leader of $\mathcal{G}_{n}$ and the leader of $\mathcal{G}_{n+1}$ execute the same algorithm, they must have the same internal state up to round $52n-$ 5, due to Theorem 3.1. In particular, they cannot give different outputs up to that round, which leads to our lower bounds on stabilization and termination for the Counting problem.
Multi-leader networks. We can now generalize Proposition 6.7 to any $\tau$ and any $\ell\geq 1$. This lower bound shows that the algorithms in Sections 4.4 and 5 are asymptotically optimal for any constant number of leaders $\ell$.
Proposition 6.8.
For any $\ell\geq 1$, no algorithm can compute the Counting function $F_{C}$ in all simple $\tau$-union-connected networks with $\ell$ leaders in less than $\tau(2n-\ell-5)$ rounds, which improves to $\tau(2n-\ell-3)$ rounds if termination is required.
Proof.
As shown in Proposition 6.7, there is a family of simple 1-union-connected networks $\mathcal{G}_{n}$, with $n\geq 1$, with the following properties. $\mathcal{G}_{n}$ has $\ell=1$ leader and $n$ processes in total; moreover, up to round $52n-$ 5, the leaders of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ have isomorphic views.
It is straightforward to see that the process $p_{\ell}$ in $\mathcal{G}^{\prime}_{n}$ and the process $p_{\ell}$ in $\mathcal{G}^{\prime}_{n+1}$, which correspond to the leaders of $\mathcal{G}_{n-\ell+1}$ and $\mathcal{G}_{n-\ell+2}$ respectively, have isomorphic views up to round $32(n-\ell)-$ 3. Since the view of $p_{1}$ is completely determined by the view of $p_{\ell}$, and it takes $\ell-1$ rounds for any information to travel from $p_{\ell}$ to $p_{1}$, we conclude that the process $p_{1}$ in $\mathcal{G}^{\prime}_{n}$ and the process $p_{1}$ in $\mathcal{G}^{\prime}_{n+1}$ have isomorphic views up to round $42n-\ell-$ 4.
It follows from Corollary 6.6 (with $k=1$, $a=2$, $b=-\ell-4$) that the Counting function with $\ell\geq 1$ leaders and $\tau=1$ cannot be computed in less than $52n-\ell-$ 5 rounds in a stabilizing fashion or in less than $32n-\ell-$ 3 if termination is required. These bounds generalize to an arbitrary $\tau$ by Proposition 2.4. ∎
Leaderless networks. Finally, we can use Proposition 6.7 to obtain a lower bound for the Average Consensus problem in leaderless networks. This lower bound shows that the leading term in the running time of the algorithm in Section 4.2, i.e., $2\tau n$, is optimal. The same holds for the algorithm in Section 4.3 assuming that $N=n$.
Proposition 6.9.
No algorithm can solve the Average Consensus problem (with or without termination) in all simple $\tau$-union-connected leaderless networks in less than $\tau(2n-6)$ rounds, which improves to $\tau(2n-4)$ rounds if termination is required.
Proof.
According to Proposition 6.7, the number of processes $n$ in a network with $\ell=1$ and $\tau=1$ cannot be determined in less than $62n-$ 6 rounds ($42n-$ 4 rounds if termination is required). We can reduce this problem to Average Consensus with $\ell=0$ and $\tau=1$ as follows. In any given network with $\ell=\tau=1$, assign input $1$ to the leader and clear its leader flag; assign input 00 to all other processes. If the processes can compute the mean input value, $1/n$, they can invert it to obtain $n$ in the same number of rounds. It follows that Average Consensus with $\ell=0$ and $\tau=1$ cannot be solved in less than $62n-$ 6 rounds ($42n-$ 4 rounds if termination is required); this immediately generalizes to an arbitrary $\tau$ by Proposition 2.4. ∎
7 Conclusions
We introduced the novel concept of history tree and used it as our main investigation technique to study computation in anonymous dynamic networks. History trees are a powerful tool that completely and naturally captures the concept of symmetry and indistinguishability among processes. We have demonstrated the effectiveness of our methods by optimally solving a wide class of fundamental problems, and we believe that our techniques will find numerous applications in other settings, as well.
We have shown that anonymous processes in $\tau$-union-connected dynamic networks can compute all the multiset-based functions and no other functions, provided that the network contains a known number of leaders $\ell\geq 1$. If there are no leaders or the number of leaders is unknown, the class of computable functions reduces to the frequency-based functions. We have also identified the Input Frequency function and the Input Multiset function as the complete problems for each class. Notably, the network’s dynamic disconnectivity $\tau$ does not affect the computability of functions, but only makes computation proportionally slower.
Moreover, we gave efficient stabilizing and terminating algorithms for computing all the aforementioned functions. Some of our algorithms make assumptions on the processes’ a-priori knowledge about the network; we proved that such assumptions are actually necessary. All our algorithms have optimal linear running times in terms of $\tau$ and the size of the network $n$.
In one case, there is still a small gap in terms of the number of leaders $\ell$. Namely, for terminating computations with $\ell\geq 1$ leaders, we have a lower bound of roughly $\tau(2n-\ell)$ rounds (Proposition 6.8) and an upper bound of roughly $\tau(\ell^{2}+\ell+1)n$ rounds (Theorem 5.2). Although these bounds asymptotically match if the number of leaders $\ell$ is constant (which is a realistic assumption in most applications), optimizing them with respect to $\ell$ is left as an open problem.
It is worth noting that for stabilizing computation (i.e., when explicit termination is not required) in networks with a constant number of leaders, our lower and upper bounds are essentially $2\tau n$ rounds; hence, in these case we were able to optimize the multiplicative constant, as well. As for terminating computation with a unique leader, we have a lower bound of $2\tau n$ rounds and an upper bound of $3\tau n$ rounds. Although we are still unable to completely close this gap, we emphasize that our findings demonstrate the practical feasibility of general computations in anonymous dynamic networks with a unique leader, which was a major open problem in this research area prior to our work.
Observe that our stabilizing algorithms use an unbounded amount of memory, as processes keep adding nodes to their views at every round. This can be avoided if the dynamic disconnectivity $\tau$ (as well as an upper bound on $n$, in the case of a leaderless network) is known: In this case, processes can run the stabilizing and the terminating version of the relevant algorithm in parallel, and stop adding nodes to their views when the terminating algorithm halts.
Our algorithms require processes to send each other explicit representations of their history trees, which have roughly cubic size in the worst case. It would be interesting to develop algorithms that only send messages of logarithmic size, possibly with a trade-off in terms of running time. We are currently able to do so for leaderless networks and networks with a unique leader, but not for networks with more than one leader [24].
Finally, we wonder if our results hold more generally for networks where communications are not necessarily synchronous. We conjecture that our algorithms can be generalized to networks where messages may be delayed by a bounded number of rounds or processes may be inactive for some rounds (provided that a “global fairness” condition is met).